# Cosmic (super)strings with a time-varying tension

### v1

@@ Appendix C Details on the dynamical system
$\frac{t_{r}}{t_{0}}\lesssim\left(\frac{\xi_{0}}{\tilde{c}_{\infty}}\frac{2(n-2)+mn-4}{n}\right)^{\frac{2n}{2(n-2)+mn-4}},$ (114)
where the constant term in (113) can be neglected, and the energy density in long strings redshifts as radiation:
$\rho=\frac{\mu(t)}{L(t)^{2}}=\frac{\mu(t_{0})}{\xi_{0}^{2}}\left(\frac{t_{0}}{t}\right)^{8/n}\sim\frac{1}{a(t)^{4}}.$ (115)
Therefore, even if the actual asymptotic solution is a genuine scaling regime, the approach to the attractor can be characterised by a long phase where the network behaves like radiation. This matches with intuitive expectations for a gas of highly relativistic strings, and can also be seen directly from (54) if $\tilde{c}\ll 1$. A similar behaviour also arises for a cosmic string network in a contracting universe Avelino:2002hx ; Avelino:2002xy , with the difference that the radiation (as opposed to scaling) regime is the actual endpoint of the evolution in that case, rather than a transient phase.
@@
$v_{r}=1\quad\quad\quad\xi_{r}=0.$ (116)
However, $\xi=0$ is actually a singular point of the system (59), and some more care should be taken. It will be convenient to reformulate the system close to the line $v=1$, and expand all coefficients to the first non-trivial order in $w\equiv 1-v$. We further assume the chopping efficiency to be given by an appropriate power of the Lorentz contraction factor $\tilde{c}(v)=\tilde{c}(1-v^{2})^{\alpha}$, with $10<\alpha<$ 1. The value $\alpha=\frac{1}{2}$ corresponds to ansatz for $\tilde{c}(v)$ motivated in Avelino:2002hx ; Avelino:2002xy . Then, (59) in the ultra-relativistic regime (with $w\ll 1$) can be recast as
$\begin{cases}\frac{{\rm d}\xi}{{\rm d}z}=-\frac{2(n-2)+mn-4}{2n}\xi+\frac{\tilde{c}}{2}(2w)^{\alpha}\\ \frac{{\rm d}w}{{\rm d}z}=-2\frac{mn-4}{n}w^{2}+k^{\prime}(1)\frac{w^{2}}{\xi},\\ \end{cases}$ (117)
@@
$\rho=\frac{w^{2-\alpha}}{\xi}\quad\quad\sigma=w^{1-\alpha},$ (118)
Eq. (117) can be further simplified to
$\begin{cases}\frac{{\rm d}\sigma}{{\rm d}z}=-2(1-\alpha)\frac{mn-4}{n}\sigma+k^{\prime}(1)\rho(1-\alpha)\\ \frac{{\rm d}\rho}{{\rm d}z}=\left(\frac{2(n-2)+mn-4}{2n}-2(2-\alpha)\frac{mn-4}{n}\right)\rho+k^{\prime}(1)(2-\alpha)\rho^{2}\sigma-\frac{\tilde{c}}{2^{1-\alpha}}\rho^{2}.\\ \end{cases}$ (119)
@@
$J_{r}=\begin{pmatrix}-2(1-\alpha)\frac{mn-4}{n}&k^{\prime}(1)(1-\alpha)\\ 0&\left(\frac{2(n-2)+mn-4}{2n}-2(2-\alpha)\frac{mn-4}{n}\right)\\ \end{pmatrix}$ (120)
@@
$\xi(t)=\xi_{0}\left(\frac{t}{t_{0}}\right)^{-\frac{1}{2n}\left[2(n-2)+mn-4\right]},$ (122)
@@
$\rho=\frac{\mu(t)}{L(t)^{2}}\sim\frac{1}{t^{8/n}}\sim\frac{1}{a(t)^{4}}.$ (123)
@@ Appendix D General momentum parameter
In the main text (and the previous appendix), we relied heavily on the explicit for of $k(v)$, which depends on a combination of analytical arguments and numerical interpolations. [^25] In particular, its functional form has been numerically verified only in the regime $v^{2}\leq\frac{1}{2}$. To make our results more robust, we can also derive bounds which hold for more arbitrary functions $k(v)$, modulo mild assumptions. Assuming a constant $\tilde{c}$, one can rewrite (56) as
$\frac{{\rm d}L}{{\rm d}t}=\frac{2L}{nt}\left[1-v^{2}\left(\frac{mn}{4}-1\right)\right]+\frac{\tilde{c}v}{2}.$ (124)
If $mn>4$ and $k(v)>0$ for $v^{2}<\frac{1}{2}$, $v^{2}>\frac{1}{2}$ asymptotically. The, (124) implies
$\frac{2L}{t}\left(\frac{2}{n}-\frac{m}{4}\right)+\frac{\tilde{c}}{2\sqrt{2}}<\frac{{\rm d}L}{{\rm d}t}<\frac{L}{t}\left(\frac{3}{n}-\frac{m}{4}\right)+\frac{\tilde{c}}{2},$ (125)
@@
$\frac{\tilde{c}t}{2\sqrt{2}\left[1-\left(\frac{4}{n}-\frac{m}{2}\right)\right]}+c_{1}t^{\frac{4}{n}-\frac{m}{2}}<L(t)-L(0)<\frac{\tilde{c}nt}{2\left[1-\left(\frac{3}{n}-\frac{m}{4}\right)\right]}+c_{2}t^{\frac{3}{n}-\frac{m}{4}}$ (126)

### v2

@@ Appendix C Details on the dynamical system
$\frac{t_{r}}{t_{0}}\lesssim\left(\frac{\xi_{0}}{\tilde{c}_{\infty}}\frac{2(n-2)+mn-4}{n}\right)^{\frac{2n}{2(n-2)+mn-4}},$ (C.14)
where the constant term in (C.13) can be neglected, and the energy density in long strings redshifts as radiation:
$\rho=\frac{\mu(t)}{L(t)^{2}}=\frac{\mu(t_{0})}{\xi_{0}^{2}}\left(\frac{t_{0}}{t}\right)^{8/n}\sim\frac{1}{a(t)^{4}}.$ (C.15)
Therefore, even if the actual asymptotic solution is a genuine scaling regime, the approach to the attractor can be characterised by a long phase where the network behaves like radiation. This matches with intuitive expectations for a gas of highly relativistic strings, and can also be seen directly from (4.16) if $\tilde{c}\ll 1$. A similar behaviour also arises for a cosmic string network in a contracting universe [98, 99], with the difference that the radiation (as opposed to scaling) regime is the actual endpoint of the evolution in that case, rather than a transient phase.
@@
$v_{r}=1\quad\quad\quad\xi_{r}=0.$ (C.16)
However, $\xi=0$ is actually a singular point of the system (4.21), and some more care should be taken. It will be convenient to reformulate the system close to the line $v=1$, and expand all coefficients to the first non-trivial order in $w\equiv 1-v$. We further assume the chopping efficiency to be given by an appropriate power of the Lorentz contraction factor $\tilde{c}(v)=\tilde{c}(1-v^{2})^{\alpha}$, with $10<\alpha<$ 1. The value $\alpha=\frac{1}{2}$ corresponds to ansatz for $\tilde{c}(v)$ motivated in [98, 99]. Then, (4.21) in the ultra-relativistic regime (with $w\ll 1$) can be recast as
$\begin{cases}\frac{{\rm d}\xi}{{\rm d}z}=-\frac{2(n-2)+mn-4}{2n}\xi+\frac{\tilde{c}}{2}(2w)^{\alpha}\\ \frac{{\rm d}w}{{\rm d}z}=-2\frac{mn-4}{n}w^{2}+k^{\prime}(1)\frac{w^{2}}{\xi},\\ \end{cases}$ (C.17)
@@
$\rho=\frac{w^{2-\alpha}}{\xi}\quad\quad\sigma=w^{1-\alpha},$ (C.18)
Eq. (C.17) can be further simplified to
$\begin{cases}\frac{{\rm d}\sigma}{{\rm d}z}=-2(1-\alpha)\frac{mn-4}{n}\sigma+k^{\prime}(1)\rho(1-\alpha)\\ \frac{{\rm d}\rho}{{\rm d}z}=\left(\frac{2(n-2)+mn-4}{2n}-2(2-\alpha)\frac{mn-4}{n}\right)\rho+k^{\prime}(1)(2-\alpha)\rho^{2}\sigma-\frac{\tilde{c}}{2^{1-\alpha}}\rho^{2}.\\ \end{cases}$ (C.19)
@@
$J_{r}=\begin{pmatrix}-2(1-\alpha)\frac{mn-4}{n}&k^{\prime}(1)(1-\alpha)\\ 0&\left(\frac{2(n-2)+mn-4}{2n}-2(2-\alpha)\frac{mn-4}{n}\right)\\ \end{pmatrix}$ (C.20)
@@
$\xi(t)=\xi_{0}\left(\frac{t}{t_{0}}\right)^{-\frac{1}{2n}\left[2(n-2)+mn-4\right]},$ (C.22)
@@
$\rho=\frac{\mu(t)}{L(t)^{2}}\sim\frac{1}{t^{8/n}}\sim\frac{1}{a(t)^{4}}.$ (C.23)
@@ Appendix D General momentum parameter
In the main text (and the previous appendix), we relied heavily on the explicit for of $k(v)$, which depends on a combination of analytical arguments and numerical interpolations. [^27] In particular, its functional form has been numerically verified only in the regime $v^{2}\leq\frac{1}{2}$. To make our results more robust, we can also derive bounds which hold for more arbitrary functions $k(v)$, modulo mild assumptions. Assuming a constant $\tilde{c}$, one can rewrite (4.18) as
$\frac{{\rm d}L}{{\rm d}t}=\frac{2L}{nt}\left[1-v^{2}\left(\frac{mn}{4}-1\right)\right]+\frac{\tilde{c}v}{2}.$ (D.1)
If $mn>4$ and $k(v)>0$ for $v^{2}<\frac{1}{2}$, $v^{2}>\frac{1}{2}$ asymptotically. The, (D.1) implies
$\frac{2L}{t}\left(\frac{2}{n}-\frac{m}{4}\right)+\frac{\tilde{c}}{2\sqrt{2}}<\frac{{\rm d}L}{{\rm d}t}<\frac{L}{t}\left(\frac{3}{n}-\frac{m}{4}\right)+\frac{\tilde{c}}{2},$ (D.2)
@@
$\frac{\tilde{c}t}{2\sqrt{2}\left[1-\left(\frac{4}{n}-\frac{m}{2}\right)\right]}+c_{1}t^{\frac{4}{n}-\frac{m}{2}}<L(t)-L(0)<\frac{\tilde{c}nt}{2\left[1-\left(\frac{3}{n}-\frac{m}{4}\right)\right]}+c_{2}t^{\frac{3}{n}-\frac{m}{4}}$ (D.3)